OpenStudy (anonymous):

describe the roots 2x^2-x-3=0

OpenStudy (anonymous):

\[\Delta = b^2 - 4ac\] \[x _{1,2} = \frac{ -b \pm \sqrt{\Delta} }{2a}\] In this case a=2 b=-1 c=-3 Do you want more help or you can pick it up from here?

OpenStudy (anonymous):

more help

OpenStudy (anonymous):

First find \[\Delta\] using the first formula.

OpenStudy (anonymous):

okay after what do you do?

OpenStudy (anonymous):

After that put \[\Delta\] in the second formula and then solve using + and - separately

OpenStudy (anonymous):

then?